# Compatibility of Kazhdan and Brauer Homomorphism

### v2

@@ Abstract.
Let $G$ be a split connected reductive group, defined over $\mathbb{Z}$. Let $F$ and $F^{\prime}$ be the non-archimedean $m$-close local fields, where $m$ is a positive integer. D.Kazhdan gave an isomorphism between the Hecke algebras ${\rm Kaz}_{m}^{F}:\mathcal{H}\big{(}G(F),K_{F}\big{)}\rightarrow\mathcal{H}\big{(}G(F^{\prime}),K_{F^{\prime}}\big{)}$, where $K_{F}$ and $K_{F^{\prime}}$ are the $m$-th usual congruence subgroups of $G(F)$ and $G(F^{\prime})$ respectively. On the other hand, if $\sigma$ is an automorphism of $G$ of prime order $l$, then we have Brauer homomorphism ${\rm Br}:\mathcal{H}(G(F),U(F))\rightarrow\mathcal{H}(G^{\sigma}(F),U^{\sigma}(F))$, where $U(F)$ and $U^{\sigma}(F)$ are compact open subgroups of $G(F)$ and $G^{\sigma}(F)$ respectively. This article aims to study the compatibility between these two maps in certain sense. Further, an application of this compatibility is given in the context of linkage–which is the representation theoretic version of Brauer homomorphism.
@@ 1. Introduction
The local Langlands correspondence relates the set of irreducible smooth representations of the group of rational points of a reductive group over a local field–with the representation theory of its Weil–Deligne group. For a split reductive group G defined over $\mathbb{Z}$, Kazhdan conjectured a link between the local Langlands correspondences for $\operatorname{{\bf G}}(F)$ and $\operatorname{{\bf G}}(F^{\prime})$, where $F$ is a $p$-adic field and $F^{\prime}$ is a non-Archimedean equal characteristic local field which are sufficiently close. Kazhdan’s approach is via isomorphisms between the respective Hecke algebras, with integral coefficients, $\mathcal{H}(\operatorname{{\bf G}}(F),K)$ and $\mathcal{H}(\operatorname{{\bf G}}(F^{\prime}),K^{\prime})$, for some specific choice of open-compact subgroups $K$ and $K^{\prime}$ of $\operatorname{{\bf G}}(F)$ and $\operatorname{{\bf G}}(F^{\prime})$ respectively (see [Kaz86, Theorem A]). Say $L$ is a non-Archimedean local field with residue characteristic $p$ and assume that $K$ is a nice compact open subgroup of $\operatorname{{\bf G}}(L)$–for instance a congruence subgroup of positive level. When $\operatorname{{\bf G}}(L)$ has an automorphism, denoted by $\sigma$, of prime order $l\neq p$ such that $\sigma(K)=K$, Treumann–Venkatesh ([TV16, Section 4]) define an $\overline{\mathbb{F}_{l}}$-algebra homomorphism between $\mathcal{H}(\operatorname{{\bf G}}(L),K)\otimes\overline{\mathbb{F}_{l}}$ and $\mathcal{H}(\operatorname{{\bf G}}^{\sigma}(L),K^{\sigma})\otimes\overline{\mathbb{F}_{l}}$, which is called the Brauer homomorphism. The Brauer homomorphism is conjectured to be compatible with Langlands functoriality principle. In this article, we study the compatibility of Brauer homomorhism and Kazhdan homomorphism in the Base change setting. Let $\mathfrak{o}_{L}$ be the ring of integers of $L$ and let $\mathfrak{p}_{L}$ be the maximal ideal of $\mathfrak{o}_{L}$. Let $E$ be a finite extension of $L$ of prime degree $l$, and let $\sigma$ be the generator of ${\rm Gal}(E/L)$. Let $K_{E}$ be the $m$-th congruence subgroup of $\operatorname{{\bf G}}(\mathfrak{o}_{E})$. Note that $K_{E}$ is stable under the action of $\sigma$ and let $K_{L}=K_{E}\cap\operatorname{{\bf G}}(L)$. Truemann and Venkatesh define Brauer homomorphism ${\rm Br}$, which is the algebra homomorphism
${\rm Br}:\mathcal{H}(\operatorname{{\bf G}}(E),K_{E})^{\sigma}\rightarrow\mathcal{H}(\operatorname{{\bf G}}(L),K_{L})$
obtained by restriction of functions on $\operatorname{{\bf G}}(E)$ to $\operatorname{{\bf G}}(L)$(see [TV16, Section 4]). The homomorphism ${\rm Br}$ is conjectured to be compatible with Langlands functoriality principle. Kazhdan ([Kaz86, Theorem A]) showed that there exists an integer $r\geq m$ such that if $F$ and $F^{\prime}$ are $r$-close (see Section (4)), there is a Hecke algebra isomorphism
${\rm Kaz}_{m}^{F}:\mathcal{H}(\operatorname{{\bf G}}(F),K_{F})\rightarrow\mathcal{H}(\operatorname{{\bf G}}(F^{\prime}),K_{F^{\prime}}).$
Kazhdan conjectured that it is sufficient to take $r=m$, and this conjecture is proved by Ganapathy in [Gan15]. It is shown in [Gan15] that Kazhdan isomorphism is compatible with the various properties in representation theory, and it also establish a bridge between local Langlands correspondence(LLC) over the field characteristic zero and that of positive characteristic.
@@
We now set up some notations to formulate the main result of this article. Let $E$ be a finite Galois extension of a non-archimedean local field $F$ of prime degree $l$. Let $F^{\prime}$ be another non-Archimedean local field which is $m$-close to $F$. Throughout this article, we assume that the prime number $l$ is different from $p$, the residue characteristic of both $F$ and $F^{\prime}$. Lemma 4.2 in Section 4 says that we have a finite Galois extension $E^{\prime}$ of $F^{\prime}$ of degree $l$ such that the fields $E$ and $E^{\prime}$ are $em$-close, where $e=1$ if $E$ is unramified and $e=l$ if $E$ is totally ramified. Let $\sigma$ (resp. $\sigma^{\prime}$) be a generator of the cyclic group ${\rm Gal}(E/F)$ (resp. ${\rm Gal}(E^{\prime}/F^{\prime})$). This leads to the following diagram:
@@
Let $F$ and $F^{\prime}$ be the non-archimedean $m$-close local fields. Let $E$ be the finite Galois extensions of $F$ of prime degree $l$, and let $E^{\prime}$ be the Galois extension of $F^{\prime}$, as indicated above. Then for any split connected reductive group G defined over $\mathbb{Z}$, we have
@@
In the context of representation theory, there is a notion of linkage (see subsection 6.2, for definition), defined by Treumann-Venkatesh–which is the representation theoretic analouge of Brauer homomorphism in the sense that they are compatible ([TV16, Section 6]). In this article, we give a precise formulation (subsection 6.3) of the compatibility of Brauer homomorphism with linkage. Then, we use this formulation and Theorem 1.1 to establish the compatibility of linkage with Kazhdan isomorphism (Theorem 6.3).
Now we briefly explain the contents of this article. In Section $2$, we review the Hecke algebra structure associated to any locally profinite group. In Section $3$, we formulate the Brauer homomorphism in our setting. The Kazdhan isomorphism and its compatibility with Galois actions will be discussed in Section $4$. In Section $5$, we prove Theorem 1.1. In Section $6$, we recall the notion of linkage and formulate some finiteness results on Tate cohomology group. Following this, we discuss the compatibility of linkage with Kazhdan isomorphism.
@@ 2. Hecke Algebra
In this section, we review the Hecke algebra structure on any locally profinite group. We refer to [TV16, Subsection 2.10] for more insight.
@@
Let $k$ be an algebraically closed field of positive characteristic $l$. Let $G$ be a locally profinite (i.e., locally compact and totally disconnected) group. Let $K$ be a compact open subgroup of $G$. We denote by $G/K$ the set of all distinct left cosets of $K$ in $G$, and it is a discrete set with a left action of $G$. For any $g\in G$, the left coset $gK$ is sometimes denoted by the symbol $[g]$. There is a natural action of $G\times G$, via left translation, on the space of all $k$-valued functions on the discrete set $G/K\times G/K$. We denote by $\mathcal{F}(G//K)$ the space of all such $k$-valued functions on $G/K\times G/K$ that are invariant under the action of the diagonal subgroup $\Delta G=\{(g,g):g\in G\}$, and whose support is a union of finitely many $G$-orbits. There is a multiplicative structure on the space $\mathcal{F}(G//K)$, given by
@@
for all $[x],[z]\in G/K$ and $f_{1},f_{2}\in\mathcal{F}(G//K)$. The space $\mathcal{F}(G//K)$ with the above multiplicative structure is a $k$-algebra, and it is called the Hecke algebra of $G$ associated with $K$. We now give an equivalent description of the Hecke algebra, which is relevant for our context.
@@
Let $\mathcal{H}(G,K)$ be the space of all $k$-valued functions which are bi-$K$-invariant i.e., $f(k_{1}gk_{2})=f(g),$ for all $k_{1},k_{2}\in K$, $g\in G$ and $f\in\mathcal{H}(G,K)$. Then the preceding bijection gives the following map
@@
where the function $\widetilde{\phi}:G\rightarrow k$ is defined as $\widetilde{\phi}(g):=\phi([1],[g])$, for all $g\in G$. The map (2.1) is clearly a bijection, and this induces a $k$-algebra structure on the space $\mathcal{H}(G,K)$. In the literature, many authors refer to the $k$-algebra $\mathcal{H}(G,K)$ as the Hecke algebra of $G$ associated with $K$.
@@ 3. Brauer Homomorphism
    There exists a subgroup $K^{\prime}$ of $K$ of finite index such that $K^{\prime}$ is an inverse limit of finite subgroups each of which is coprime to $l$.
@@
for all $x,y\in G$ and $\phi\in\mathcal{F}(G//K)$. It is shown in [TV16, Section 4] that the Brauer homomorphism ${\rm Br}$ is a $k$-algebra homomorphism from $\mathcal{F}(G//K)^{\sigma}$ to $\mathcal{F}(G^{\sigma}//K^{\sigma})$, and it is the restriction map
@@
for all $g\in G$. Then we have a natural induced action of $\langle\sigma\rangle$ on the space $\mathcal{H}(G,K)$, defined as
@@
for all $g\in G$. Note that the bijection (2.1) preserves the action of the group $\langle\sigma\rangle$. Then we have commutative diagram
@@
where the horizontal arrows in the diagram are the bijections induced by (2.1). We observe that the $k$-algebra map ${\rm\overline{Br}}:\mathcal{H}(G,K)^{\sigma}\rightarrow\mathcal{H}(G^{\sigma},K^{\sigma})$, induced by the above diagram, is the restriction map to subgroup $G^{\sigma}$ i.e.,
@@ 4. Kazhdan isomorphism
In this section, we review Kazhdan isomorphism and set up some initial results. Let $F$ and $F^{\prime}$ be the non-Archimedean local fields, both of residue characteristic $p$, with ring of integers $\mathfrak{o}_{F}$ and $\mathfrak{o}_{F^{\prime}}$ and its maximal ideals $\mathfrak{p}_{F}$ and $\mathfrak{p}_{F^{\prime}}$ respectively. The fields $F$ and $F^{\prime}$ are called $m$-close if there is a ring isomorphism $\Lambda:\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m}\rightarrow\mathfrak{o}_{F^{\prime}}/\mathfrak{p}_{F^{\prime}}^{m}$. We choose uniformizer $\varpi$ (resp. $\varpi^{\prime}$) of the field $F$ (resp. $F^{\prime}$) such that the class of $\varpi$ corresponds to the class of $\varpi^{\prime}$ via the map $\Lambda$.
@@
Let G be a split connected reductive group defined over $\mathbb{Z}$. Fix a Borel subgroup B of G, defined over $\mathbb{Z}$. Let $\textbf{B}=\textbf{T}\textbf{U}$ be a Levi decomposition over $\mathbb{Z}$ with maximal torus T and unipotent radical U. We denote by $X^{*}(\textbf{T})$ and $X_{*}(\textbf{T})$ the character lattice and the cocharacter lattice of T respectively. Then there is a natural $\mathbb{Z}$-bilinear pairing
@@
where $(\lambda\circ\alpha)(x)=x^{\langle\alpha,\lambda\rangle}$, for all $x\in\mathbb{G}_{m}$. Let $\Phi\subseteq X^{*}(\textbf{T})$ be the set of roots of T in G, and let $\Phi^{+}$ be the set of positive roots of T in B. Let $G_{F}=\textbf{G}(F)$. Consider the subgroup
@@
Then $K_{F}$ is a compact open subgroup of $G_{F}$. Fix a Haar measure $\mu_{F}$ on $G_{F}$ such that $\mu_{F}(K_{F})=1$. For $g\in G_{F}$, let $t_{g}=1_{K_{F}gK_{F}}$, the characteristic function on the double coset $K_{F}gK_{F}$. Then the Hecke algebra $\mathcal{H}(G_{F},K_{F})$ is generated as a $k$-vector space by $\{t_{g}:g\in G_{F}\}$. We denote by $\mathcal{H}_{\mathbb{Z}}(G_{F},K_{F})$ the space generated by $\{t_{g}:g\in G_{F}\}$ over $\mathbb{Z}$. Set
@@
for all $x,y\in\textbf{G}(\mathfrak{o}_{F})$ and $g\in G_{F}(\mu)$. Then the finite group $H_{m}:=\textbf{G}(\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m})\times\textbf{G}(\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m})$ acts transitively on the set $\{K_{F}xK_{F}:x\in G_{F}(\mu)\}$. Let $\Gamma_{\mu}(F)$ be the stabilizer of $K_{F}\varpi_{\mu}K_{F}$ in $H_{m}$. Let $T_{\mu}(F)\subseteq\textbf{G}(\mathfrak{o}_{F})\times\textbf{G}(\mathfrak{o}_{F})$ be the set of representatives of the coset space $H_{m}/\Gamma_{\mu}(F)$ under the composition
@@
For each $\mu\in X_{*}(\textbf{T})^{-}$, the map $\Lambda$ induces a bijection $T_{\mu}(F)\rightarrow T_{\mu}(F^{\prime})$. Under this bijection, there is an isomorphism ([Kaz86, Section 1]) of $\mathbb{Z}$-modules:

### v4

@@ Abstract.
Let $G$ be a connected split reductive group defined over $\mathbb{Z}$. Let $F$ and $F^{\prime}$ be two non-Archimedean $m$-close local fields, where $m$ is a positive integer. D.Kazhdan gave an isomorphism between the Hecke algebras ${\rm Kaz}_{m}^{F}:\mathcal{H}\big{(}G(F),K_{F}\big{)}\rightarrow\mathcal{H}\big{(}G(F^{\prime}),K_{F^{\prime}}\big{)}$, where $K_{F}$ and $K_{F^{\prime}}$ are the $m$-th usual congruence subgroups of $G(F)$ and $G(F^{\prime})$ respectively. On the other hand, if $\sigma$ is an automorphism of $G$ of prime order $l$, then we have Brauer homomorphism ${\rm Br}:\mathcal{H}(G(F),U(F))\rightarrow\mathcal{H}(G^{\sigma}(F),U^{\sigma}(F))$, where $U(F)$ and $U^{\sigma}(F)$ are compact open subgroups of $G(F)$ and $G^{\sigma}(F)$ respectively. In this article, we study the compatibility between these two maps in the local base change setting. Further, an application of this compatibility is given in the context of linkage–which is the representation theoretic version of Brauer homomorphism.
@@ 1. Introduction
The local Langlands correspondence relates the set of irreducible smooth complex representations of the group of rational points of a reductive group over a local field–with the representation theory of its Weil–Deligne group. For a connected split reductive group G defined over $\mathbb{Z}$, D. Kazhdan conjectured a link between the local Langlands correspondences for $\operatorname{{\bf G}}(F)$ and $\operatorname{{\bf G}}(F^{\prime})$, where $F$ is a $p$-adic field and $F^{\prime}$ is a non-Archimedean positive characteristic local field–which are sufficiently close. Kazhdan’s approach is via isomorphisms between the respective Hecke algebras with complex coefficients, $\mathcal{H}(\operatorname{{\bf G}}(F),K)$ and $\mathcal{H}(\operatorname{{\bf G}}(F^{\prime}),K^{\prime})$, for some specific choice of open-compact subgroups $K$ and $K^{\prime}$ of $\operatorname{{\bf G}}(F)$ and $\operatorname{{\bf G}}(F^{\prime})$ respectively (see [Kaz86, Theorem A]). Say $L$ is a non-Archimedean local field with residue characteristic $p$ and assume that $K$ is a nice compact open subgroup of $\operatorname{{\bf G}}(L)$–for instance a congruence subgroup of positive level. When $\operatorname{{\bf G}}(L)$ has an automorphism, denoted by $\sigma$, of prime order $l\neq p$ such that $\sigma(K)=K$, Treumann–Venkatesh ([TV16, Section 4]) define an $\overline{\mathbb{F}}_{l}$-algebra morphism between Hecke algebras with $\overline{\mathbb{F}}_{l}$-coefficients, $\mathcal{H}(\operatorname{{\bf G}}(L),K)^{\sigma}$ and $\mathcal{H}(\operatorname{{\bf G}}^{\sigma}(L),K^{\sigma})$, which is called the Brauer homomorphism, and it is conjectured to be compatible with Langlands functoriality. This article aims to study the compatibility between the Brauer homomorphism and Kazhdan isomorphism in the local base change setting. So far, this type of compatibility has not been addressed in the literature. Using this compatibility, one can relate the functoriality principles for connected reductive algebraic groups defined over the close local fields $F$ and $F^{\prime}$. This article gives one such treatment in the context of linkage (see [TV16, Section 6])–which is related to the functoriality over $\overline{\mathbb{F}}_{l}$.
Let $\mathfrak{o}_{F}$ be the ring of integers of $F$, and let $\mathfrak{p}_{F}$ be the maximal ideal of $\mathfrak{o}_{F}$. Let $E$ be a Galois extension of $F$ of prime degree $l$, where $l$ is different from the residue characteristic of $F$. Let $\sigma$ be a generator of ${\rm Gal}(E/F)$. Similar notations follow for $E$. Let $K_{E}$ (resp. $K_{F}$) be the $m$-th congruence subgroup of $\operatorname{{\bf G}}(\mathfrak{o}_{E})$ (resp. $\operatorname{{\bf G}}(\mathfrak{o}_{F})$). Note that $K_{E}$ is stable under the action of $\sigma$ and let $K_{F}=K_{E}\cap\operatorname{{\bf G}}(F)$. Treumann–Venkatesh ( [TV16, Section 4]) define Brauer homomorphism, denoted by ${\rm Br}$–which is the $\overline{\mathbb{F}}_{l}$-algebra homomorphism
${\rm Br}:\mathcal{H}(\operatorname{{\bf G}}(E),K_{E})^{\sigma}\rightarrow\mathcal{H}(\operatorname{{\bf G}}(F),K_{F}),$
obtained by restriction of functions on $\operatorname{{\bf G}}(E)$ to $\operatorname{{\bf G}}(F)$. On the other hand, following Kazhdan’s approach in [Kaz86, Theorem A], we get an $\overline{\mathbb{F}}_{l}$-algebra isomorphism (Proposition 4.5)
${\rm Kaz}_{m}^{F}:\mathcal{H}(\operatorname{{\bf G}}(F),K_{F})\rightarrow\mathcal{H}(\operatorname{{\bf G}}(F^{\prime}),K_{F^{\prime}}),$
provided the fields $F$ and $F^{\prime}$ are sufficiently close. The above isomorphism ${\rm Kaz}_{m}^{F}$ was originally constructed for complex Hecke algebras (Proposition 4.1)–which relies on the finiteness of complex Hecke algebras, due to Bernstein ([Ber84, Corollary 3.4]). For arbitrary Noetherian $\mathbb{Z}_{l}$-algebra $R$, where $\mathbb{Z}_{l}$ is the ring of integers of the field of $l$-adic numbers, the finiteness of Hecke algebras with coefficients in $R$ follows from the work of [DHKM24, Theorem 1.1]. It enables us to define the isomorphism ${\rm Kaz}_{m}^{F}$ between mod-$l$ Hecke algebras.
The method of close local fields is fruitful in obtaining analogous results for reductive groups defined over local fields of positive characteristic–which are known for reductive groups over $p$-adic fields. Using the close local fields approach, A.I. Badulescu ([Bad02]) proved the Jacquet-Langlands correspondence for the general linear group ${\rm GL}_{n}(F^{\prime})$. A similar technique was used by Badulescu–Henniart–Lemaire–Sécherre ([BHLS10]) to classify the smooth unitary dual of ${\rm GL}_{n}(D)$, where $D$ is a central division algebra over $F^{\prime}$. Kazhdan isomorphism and its various properties have been further studied by several others, notably by R. Ganapathy (see [Gan15], [Gan22]).
@@
We now state the main results of this article. Let G be a connected split reductive group defined over $\mathbb{Z}$. Let $F$ and $F^{\prime}$ be two non-Archimedean local fields with the same residue characteristic $p$. We assume that $F$ and $F^{\prime}$ are $m$-close (i.e., there exists a ring isomorphism $\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m}\xrightarrow{\sim}\mathfrak{o}_{F^{\prime}}/\mathfrak{p}_{F^{\prime}}^{m}$). Let $E/F$ be a finite Galois extension of prime degree $l$ with $l\neq p$. Then there exists a finite Galois extension $E^{\prime}/F^{\prime}$ of degree $l$ such that $E^{\prime}$ is $em$-close to $E$, where $e=1$ if $E/F$ is unramified, and $e=l$ if $E/F$ is totally ramified (Lemma 4.6). Let $\sigma$ (resp. $\sigma^{\prime}$) be a generator of the group ${\rm Gal}(E/F)$ (resp. ${\rm Gal}(E^{\prime}/F^{\prime})$). Then we have the Kazhdan isomorphism ${\rm Kaz}_{em}^{E}$ between the mod-$l$ Hecke algebras $\mathcal{H}(\textbf{G}(E),K_{E})$ and $\mathcal{H}(\textbf{G}(E^{\prime}),K_{E^{\prime}})$, and it is equivariant under the Galois actions (Proposition 4.9)–which leads to the following diagram:
@@
Let $F$ and $F^{\prime}$ be two non-Archimedean $m$-close local fields with residue characteristic $p$. Let $E$ be a finite Galois extension of $F$ of prime degree $l$ with $l\neq p$, and let $E^{\prime}$ be the Galois extension of $F^{\prime}$, as indicated above. Then, for any connected split reductive group G defined over $\mathbb{Z}$, we have
@@
Using the above theorem, we prove the compatibility of linkage with Kazhdan isomorphism. Say $(\pi,V)$ is an irreducible smooth $\overline{\mathbb{F}}_{l}$-representation of $\textbf{G}(E)$ such that $\pi$ is isomorphic to twisted representation $\pi^{\sigma}$. Then we have the Tate cohomology groups $\widehat{H}^{i}(\sigma,V)$, for $i\in\{0,1\}$, defined as $\overline{\mathbb{F}}_{l}$-representations of $\textbf{G}(F)$. An irreducible sub–quotient $\rho$ of $\widehat{H}^{i}(\sigma,V)$ is said to be linked with $\pi$. Let $\rho^{\prime}$ (resp. $\pi^{\prime}$) be the smooth irreducible $\overline{\mathbb{F}}_{l}$-representation of $\textbf{G}(F^{\prime})$ (resp. $\textbf{G}(E^{\prime})$) associated with $\rho$ (resp. $\pi$) via Kazhdan isomorphism and the natural correspondence between the set of simple modules over Hecke algebras and the set of irreducible representations of locally profinite group (see [Gan15, Section 2.3]). Then, using Theorem 1.1 and the compatibility of Brauer homomorphism with linkage (see subsection 6.4), we show that $\rho^{\prime}$ is linked with $\pi^{\prime}$ if and only if $\rho$ is linked with $\pi$ (Theorem 6.5).
We now briefly explain the contents of this article. Section $2$ reviews the Hecke algebra structure associated with any locally profinite group. In Section $3$, we discuss the Brauer homomorphism. In Section $4$, we review the Kazdhan isomorphism and formulate its mod-$l$ version. We also set up some initial results, which are crucial for the main result. Theorem 1.1 is proved in Section $5$. In Section $6$, we prove the compatibility of linkage with the Kazhdan isomorphism. It also includes some finiteness results on the Tate cohomology groups.
@@ 2. Hecke Algebra
This section reviews the Hecke algebra structure on any locally profinite group. We refer to [TV16, Subsection 2.10] for more insight.
@@
Let $k$ be an algebraically closed field of positive characteristic $l$. Let $G$ be a locally profinite (i.e., locally compact and totally disconnected) group. Let $K$ be a compact open subgroup of $G$. We denote by $G/K$ the set of all distinct left cosets of $K$ in $G$, and it is a discrete set with a left action of $G$. For any $g\in G$, we sometimes use the notation $[g]$ to denote the left coset $gK$. There is a natural action of $G\times G$, via left translation, on the space of $k$-valued functions on the discrete set $G/K\times G/K$. We denote by $\mathcal{F}(G//K)$ the space of all such $k$-valued functions on $G/K\times G/K$ that are invariant under the action of the diagonal subgroup $\Delta G=\{(g,g):g\in G\}$, and whose support is a union of finitely many $G$-orbits. There is a multiplicative structure in the space $\mathcal{F}(G//K)$, given by
@@
for all $[x],[z]\in G/K$ and $f_{1},f_{2}\in\mathcal{F}(G//K)$. The space $\mathcal{F}(G//K)$ with the above multiplicative structure is a $k$-algebra, and it is called the Hecke algebra of $G$ associated with $K$. We now give an equivalent description of the Hecke algebra relevant to our context.
@@
Let $\mathcal{H}(G,K)$ be the space of all $k$-valued functions which are bi-$K$-invariant i.e., $f(k_{1}gk_{2})=f(g),$ for all $k_{1},k_{2}\in K$, $g\in G$ and $f\in\mathcal{H}(G,K)$. Then, the preceding bijection gives the following map
@@
where the function $\widetilde{\phi}:G\rightarrow k$ is defined as $\widetilde{\phi}(g):=\phi([1],[g])$, for all $g\in G$. The map (2.1) is a bijection, and it induces a $k$-algebra structure on the space $\mathcal{H}(G,K)$. In the literature, many authors refer to the $k$-algebra $\mathcal{H}(G,K)$ as the Hecke algebra of $G$ associated with $K$.
@@ 3. Brauer Homomorphism
    There exists a subgroup $K^{\prime}$ of $K$ of finite index such that $K^{\prime}$ is an inverse limit of finite subgroups, each of which is coprime to $l$.
@@
for all $x,y\in G$ and $\phi\in\mathcal{F}(G//K)$. In the article [TV16, Section 4], the authors proved that the Brauer homomorphism ${\rm Br}$ is a $k$-algebra homomorphism from $\mathcal{F}(G//K)^{\sigma}$ to $\mathcal{F}(G^{\sigma}//K^{\sigma})$, and defined by the restriction
@@
for all $g\in G$. Then, we have a natural induced action of $\langle\sigma\rangle$ on the space $\mathcal{H}(G,K)$, defined as
@@
for all $g\in G$. Note that the bijection (2.1) preserves the action of the group $\langle\sigma\rangle$. Then, we have a commutative diagram
@@
Here, the horizontal arrows in the diagram are the bijections induced by (2.1). We observe that the $k$-algebra map ${\rm\overline{Br}}:\mathcal{H}(G,K)^{\sigma}\rightarrow\mathcal{H}(G^{\sigma},K^{\sigma})$, induced by the above diagram, is the restriction map to subgroup $G^{\sigma}$, i.e.,
@@ 4. Kazhdan isomorphism
In this part, we review Kazhdan isomorphism and set up some initial results. Let $F$ and $F^{\prime}$ be the non-Archimedean local fields, both of residue characteristic $p$, with ring of integers $\mathfrak{o}_{F}$ and $\mathfrak{o}_{F^{\prime}}$ and its maximal ideals $\mathfrak{p}_{F}$ and $\mathfrak{p}_{F^{\prime}}$ respectively. The fields $F$ and $F^{\prime}$ are called $m$-close if there is a ring isomorphism $\Lambda:\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m}\rightarrow\mathfrak{o}_{F^{\prime}}/\mathfrak{p}_{F^{\prime}}^{m}$. We choose uniformizer $\varpi$ (resp. $\varpi^{\prime}$) of the field $F$ (resp. $F^{\prime}$) such that the class of $\varpi$ corresponds to the class of $\varpi^{\prime}$ via the map $\Lambda$.
@@
Let G be a split connected reductive group defined over $\mathbb{Z}$. Fix a Borel subgroup B of G, defined over $\mathbb{Z}$. Let $\textbf{B}=\textbf{T}\textbf{U}$ be a Levi decomposition over $\mathbb{Z}$ with maximal torus T and unipotent radical U. We denote by $X^{*}(\textbf{T})$ and $X_{*}(\textbf{T})$ the character lattice and the co-character lattice of T respectively. Then, there is a natural $\mathbb{Z}$-bilinear pairing
@@
where $(\lambda\circ\alpha)(x)=x^{\langle\alpha,\lambda\rangle}$, for all $x\in\mathbb{G}_{m}$. Let $\Phi\subseteq X^{*}(\textbf{T})$ be the set of roots of T in G, and let $\Phi^{+}$ be the set of positive roots of T in B.
4.1.1.
Let $G_{F}$ denotes the group $\textbf{G}(F)$. Consider the subgroup
@@
Then $K_{F}$ is a compact open subgroup of $G_{F}$. Fix a Haar measure $\mu_{F}$ on $G_{F}$ such that $\mu_{F}(K_{F})=1$. For each $g\in G_{F}$, we denote by $t_{g}$ the characteristic function on the double coset $K_{F}gK_{F}$. Then the Hecke algebra $\mathcal{H}(G_{F},K_{F})$ is generated as a $k$-vector space by $\{t_{g}:g\in G_{F}\}$. We denote by $\mathcal{H}_{\mathbb{Z}}(G_{F},K_{F})$ the space generated by $\{t_{g}:g\in G_{F}\}$ over $\mathbb{Z}$. Set
@@
for all $x,y\in\textbf{G}(\mathfrak{o}_{F})$ and $g\in G_{F}(\mu)$. Then, the finite group $H_{m}:=\textbf{G}(\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m})\times\textbf{G}(\mathfrak{o}_{F}/\mathfrak{p}_{F}^{m})$ acts transitively on the set $\{K_{F}xK_{F}:x\in G_{F}(\mu)\}$. Let $\Gamma_{\mu}(F)$ be the stabilizer of $K_{F}\varpi_{\mu}K_{F}$ in $H_{m}$. Let $T_{\mu}(F)\subseteq\textbf{G}(\mathfrak{o}_{F})\times\textbf{G}(\mathfrak{o}_{F})$ be the set of representatives of the coset space $H_{m}/\Gamma_{\mu}(F)$ under the composition
@@
4.1.2.
For each $\mu\in X_{*}(\textbf{T})^{-}$, the isomorphism $\Lambda$ induces a bijection $T_{\mu}(F)\rightarrow T_{\mu}(F^{\prime})$. Following Kazhdan’s constructions in [Kaz86, Section 1], we get an isomorphism of $\mathbb{Z}$-modules:
$\mathcal{H}_{\mathbb{Z}}(G_{F},K_{F})\longrightarrow\mathcal{H}_{\mathbb{Z}}(G_{F^{\prime}},K_{F^{\prime}})$ (4.3)
and this induces a $\mathbb{C}$-vector space isomorphism
${\rm Kaz}_{m,\mathbb{C}}^{F}:\mathcal{H}_{\mathbb{C}}(G_{F},K_{F})\longrightarrow\mathcal{H}_{\mathbb{C}}(G_{F^{\prime}},K_{F^{\prime}}),$
where $\lambda\in X_{*}(\textbf{T})^{-}$, $(a_{i},a_{j})\in T_{\lambda}(F)$ and $(a_{i}^{\prime},a_{j}^{\prime})$ is the corresponding element in $T_{\lambda}(F^{\prime})$. The following result is due to Kazhdan ([Kaz86, Theorem A]).
Proposition 4.1.
Given $m\geq 1$. There exists a positive integer $r\geq m$ such that if $F$ and $F^{\prime}$ are $r$-close, the map ${\rm Kaz}_{m,\mathbb{C}}^{F}$, defined above, is infact an isomorphism of $\mathbb{C}$-algebras.
In order to establish that the vector space isomorphism ${\rm Kaz}_{m,\mathbb{C}}^{F}$ is compatible with algebra structures, the fields need to be a few levels closer. Kazhdan conjectured that one could take $r=m$ in Proposition 4.1–which was proved by R. Ganapathy ([Gan15, Corollary 3.15]).
4.2. Kazhdan isomorphism over positive characteristic
In this subsection, we formulate a version of Proposition 4.1 for mod-$l$ Hecke algebras. Recall that $\mathcal{H}(G_{F},K_{F})$ denotes the Hecke algebra, consisting of all $k$-valued smooth, compactly supported and bi-$K_{F}$-invariant functions on $G_{F}$. Suppose $F$ and $F^{\prime}$ are $m$-close. We denote by $X_{m}(F)$ the set of all $K_{F}$ double cosets in $G_{F}$. Similar notations are followed for $F^{\prime}$. The bijection $\phi:X_{m}(F)\rightarrow X_{m}(F^{\prime})$ gives the $\mathbb{Z}$-module isomorphism (see [Kaz86, Section 1]) of integral Hecke algebras